OpenStudy (anonymous):

Find the solution to the system using a table. Write your answers as ordered pairs using parentheses. If there are two solutions, separate the two solutions with a comma. Do not put a space between them. y = 2x - 7 y = -5x

OpenStudy (texaschic101):

y = -5x now sub -5x in for y in the other equation -5x = 2x - 7 -5x - 2x = - 7 -7x = -7 x = 1 y = -5x y = -5(1) y = -5 so when x = 1, y = -5....(1,-5) when x = 2, y = -10 .....(2,-10) when x = 3, y = -15 .....(3,-15) when x = 4, y = -20......(4,-20) when x = 5, y = - 25....(5,-25)

OpenStudy (anonymous):

Thank you